Remember Little Johnny? He now has two investments that produce a $150 income each month. If $1,000 more is invested at 9% than at 10%, how much was invested at each percent?

Answer by mathtech5(4)

Let x represent the amount invested in the 10%. Then (x+100) would be the amount invested at 9%.
Set up the equation: .1x + .09(x+1000)= 150
.1x + .09x + 100 = 150
.19x + 100 = 150 (subtract 100 from both sides)
.19x = 50 (divide by .19)
x = 263.16
Therefore, the amount invested at 10% was $263.16 and the amount invested in 9% was $1263.16.